OpenStudy (anonymous):

Which of the following represent the zeros of f(x) = 6x^3 – 31x^2 + 4x + 5 ?

OpenStudy (anonymous):

5, –one–third, one–half 5, one–third, – one–half –5, one–third, one–half 5, one–third, one–half

OpenStudy (diyadiya):

(x-5)(3x+1)(2x-1)=0 x=5 x=-1/3 x=1/2 5, –one–third, one–half